OpenStudy (anonymous):

how to add -x^3/5+ 1/5x^2/5

OpenStudy (shadowfiend):

Wait so is that: \[\frac{-x^3}{5} + \frac{\frac{1}{5x^2}}{5}\] ?

OpenStudy (anonymous):

-x^(3/5) + 1/5x^(2/5)